OpenStudy (anonymous):

Pls give steps. Most probably to be solved using simultaneous linear equations (i.e. pair of linear equations in two variables) Father's age is 3 times the sum of ages of his two children. After 5 years his age will be twice the sum of ages of two children. Find the age of father. Father's age

OpenStudy (lalaly):

let the sum of ages of children =x fathers age =y Fathers age is 3 times the sum of ages of his two chidren: y=3x After 5 years his age will be twice the sum of ages of two children y+5=2(x+5) now solve

OpenStudy (anonymous):

@lalaly substituting y = 3x 3x + 5 = 2x + 10 3x - 2x = 10 - 5 x = 5 so y = 3x = 3*5 = 15 A father of 15 years age !!!!! I doubt it....☺

OpenStudy (anonymous):

I am getting answer as 45 years..

OpenStudy (lalaly):

how did u do it?

OpenStudy (anonymous):

Let age of one child be x years and age of second child be y years then father's age will be 3(x + y) years i.e 3x + 3y After 5 years first child is x+5 years second child y+5 years father's age will be 3x + 3y + 5 ATQ, 3x + 3y + 5 = 2(x+5 + y+5) 3x + 3y + 5 = 2x + 2y + 20 x + y = 15 so fathers age i.e 3(x+y) = 3(15) = 45 Correct??????

OpenStudy (lalaly):

yeh ur right my bad=)

OpenStudy (anonymous):

No problem. We all go zonkers sometimes....☺☻ I was not sure that is why I posted it here......

OpenStudy (lalaly):

lol goodjob;)

OpenStudy (anonymous):

Still, thanks for trying. Can u look-up my other question on pair of linear equations pls...

OpenStudy (lalaly):

Sure:D